Dorthy is 14 years youngfer than Rita. Ten years ago Rita was 3 times as old as Dorthy was then. How old is each now?
Answer by Maths68(1474)

Let
Present age of Dorthy = d
Present age of Rita = r

Dorthy is 14 years younger than Rita.
d=r-14.........(1)
Ten years ago
Dorthy age was = d-10
Rita age was = r-10
Then,
Rita was 3 times as old as Dorthy
r-10=3(d-10)
r-10=3d-30
r=3d-30+10
r=3d-20...........(2)
Plug in the value of d from (1) to (2)
r=3(r-14)-20
r=3r-42-20
r-3r=-62
-2r=-62
-2r/-2=-62/-2
r=31
Pulg in the value of r in (1)
d=r-14
d=31-14
d=17


Present age of Dorthy = d = 17 years old
Present age of Rita = r = 31 years old